# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ 3.3 Minimax Regret Bounds
It should be noted that by strong convexity, this implies a lower bound on the tracking error (up to constant factors). Furthermore, it is worth noting that under uniformly spread drift (i.e., $q=1$, $p=\infty$, and $\Delta\asymp\mathcal{V}_{T}/(\mu T)$), optimizing the constant stepsize in Theorem˜3.4 yields a dynamic regret upper bound of the form $\mathcal{R}_{T}\;\lesssim\;(1-\beta)^{-2/3}\,\sigma^{4/3}\,\mu^{-1/3}\,\mathcal{V}_{T}^{2/3}\,T^{1/3}$ which matches the corresponding minimax lower bound in its dependence on $\mathcal{V}_{T}$, $T$, $\sigma$ and crucially the momentum penalty $(1-\beta)^{-2/3}$.
The minimax regret bound consists of two terms. The first term corresponds to a statistical regime in which regret is forced primarily by noisy gradient feedback together with the gradient-variation budget. In this regime, performance is bottlenecked by information rather than by the algorithm’s internal dynamics: no method can extract substantially more signal from the same noisy gradients under the same variation constraint. For SGDM, the dependence on $(1-\beta)$ in this term is comparatively mild (it enters through the stability-restricted tuning of the constant stepsize), reflecting that this regime is fundamentally governed by noise and variation rather than inertia. The second term is induced by a inertia-limited regime where delay dominates. Momentum reduces the impact of stochastic noise by averaging gradients over time, but this same averaging creates inertia: after a change in the environment, the iterate continues to follow stale gradient information. Thus, even under the best allowable tuning, high momentum creates an unavoidable "inertia window" of length $\Omega(\kappa/(1-\beta))$ where $\kappa=L/\mu$ under standard stability step-size restrictions.
These two terms have two concrete implications. First, in our upper bounds the drift-driven component of tracking error/regret is explicitly amplified as $\beta\uparrow 1$, whereas plain SGD does not incur this inertia amplification. Second, this phenomenon is not an artifact of analysis: the minimax lower bound constructs blockwise shifts (consistent with the gradient-variation budget) for which any $\mathrm{SGDM}(\beta)$ policy must spend $\tau_{\beta}$ steps per change in a transient regime where it is systematically misaligned with the current optimizer, thereby forcing an unavoidable regret penalty that deteriorates with $\beta$. Taken together, the theory explains the empirical dichotomy: in near-stationary regimes the noise term dominates and momentum can help, while under genuine distribution shift the inertia term dominates and SGD can be uniformly preferable because it adapts immediately rather than averaging gradients over time.
Figure 2: Two minimax regimes and the inertia window. (a) After a regime change (distribution shift), momentum averages past gradients: this reduces noise but induces inertia, so the iterate lags behind the drifting minimizer for an “inertia window” whose length grows with the momentum parameter. (b) The minimax lower bound in Theorem˜3.7 is the maximum of two contributions: a noise/variation-limited term which is bottlenecked by information rather than inertia, and an inertia-limited term that worsens with momentum and dominates when delay is the bottleneck. The crossover separates a statistical regime from an inertia regime, explaining when momentum is provably worse than SGD.
4 Experimental Results
In this section, we empirically corroborate our theoretical findings by evaluating the tracking performance of SGD and momentum-based methods across a hierarchy of optimization tasks. We study online optimization under drifting optima by generating a stream of losses whose population minimizer $\bm{\theta}_{t}^{\star}$ varies over time. We begin with a strongly convex quadratic objective, providing a direct bridge to our theoretical bounds where the impact of drift and noise can be isolated. We then consider three more advanced tasks: (i) linear regression (squared loss) with a drifting ground-truth weight vector $\bm{\theta}_{t}^{\star}$, (ii) logistic regression (binary cross-entropy) with drifting $\bm{\theta}_{t}^{\star}$, and (iii) a teacher–student two-layer MLP regression problem (MSE) with drifting teacher parameters. Across all settings, we simulate non-stationarity via a random walk of the target parameters. We compare SGD, Heavy-Ball (HB), and Nesterov accelerated gradient (NAG). For our last three experiments, we use a common base momentum $\beta=0.9$ for HB/NAG and task/regime-specific base step sizes $\gamma$ (reported in Table˜2). Crucially, to reflect non-stationary stability constraints for momentum, we cap momentum step sizes by the theoretical bound
$\gamma\;\leq\;\frac{\mu(1-\beta)^{2}}{4L^{2}},$
which can be orders of magnitude smaller than the SGD stability range (e.g., $\gamma_{\max}^{\mathrm{mom}}=2.5\times 10^{-5}$ for $\kappa=10$ and $2.5\times 10^{-9}$ for $\kappa=1000$ when $\beta=0.9$). This cap isolates the empirically observed sluggishness of momentum under drift: momentum must use a tiny stable step size and therefore adapts slowly to distribution shifts.
4.1 Strongly Convex Quadratic Minimizer
At each time step $t\geq 0$, we consider the time-varying objective function
$f_{t}(\bm{\theta})=\frac{\mu}{2}\|\bm{\theta}-\bm{\theta}_{t}^{\star}\|^{2},$ (4.1)
where $\bm{\theta}\in\mathbb{R}^{d}$ is the parameter vector and $\bm{\theta}_{t}^{\star}\in\mathbb{R}^{d}$ denotes the (moving) minimizer at time $t$. We set $\mu=1$. The function is $\mu$-strongly convex with gradient $\nabla f_{t}(\bm{\theta})=\mu(\bm{\theta}-\bm{\theta}_{t}^{\star}).$ The sequence of optima $\{\bm{\theta}_{t}^{\star}\}_{t\geq 0}$ is modeled as a random walk in $\mathbb{R}^{d}$:

### v3

@@ E.5 Obtaining a minimax lower bound in the statistical/variation-limited regime
Proof of Lemma˜E.13.
First note that by Corollary˜E.2, we have that $\mathbb{P}(\widehat{U}\neq U)\geq\frac{1}{2}$. Subsequently invoking Lemma˜E.1, we conclude that there does not exist an admissible policy $\pi\in\Pi_{\beta}$ such that $\sup_{g\in\mathcal{G}_{p,q}(\mathbb{V}_{T})}\mathcal{R}_{T}^{\pi,\phi^{G}}(g)\;\leq\;\frac{1}{9}\,\inf_{\begin{subarray}{c}g,\tilde{g}\in\Theta\end{subarray}}\sum_{t=1}^{T}\chi(g_{t},\tilde{g}_{t})$. This gives us a lower bound for the minimax regret. Now we must just enforce the gradient variational budget and the information constraint. To enforce the gradient variational budget, we have by Lemma˜E.10 that the condition $\mathrm{GVar}_{p,q}(G^{u}_{1:T})\leq\mathbb{V}_{T}$ for all $u\in\mathcal{U}$ is ensured whenever $C\mu a^{\alpha}J^{1/q}\leq\mathbb{V}_{T}$, which is exactly achieved by the choice (209).
To enforce the information (Fano/KL) constraint,, we have by Lemma˜E.12, for any policy $\pi$ and any $u,v\in\mathcal{U}$,
Combining this uniform pairwise bound with the standard mutual-information upper bound (Lemma˜G.7), we get
$I(U;Z_{1:T})\ \leq\ C\,\frac{\mu^{2}a^{2}}{\sigma^{2}/(1-\beta)}\,T.$
Since $\log|\mathcal{U}|\geq cJ$, imposing (210) makes $I(U;Z_{1:T})$ a small constant fraction of $\log|\mathcal{U}|$. Fano’s inequality then yields a constant lower bound on the minimax probability of misidentifying $U$, uniformly over all estimators $\widehat{U}(Z_{1:T})$.
Substituting (209) into (210) gives
$C\,\frac{\mu^{2}T}{\sigma^{2}/(1-\beta)}\left(\frac{\mathbb{V}_{T}}{C\mu J^{1/q}}\right)^{\!2/\alpha}\ \lesssim\ J,$
or equivalently,
$\frac{\mu^{2-2/\alpha}}{\sigma^{2}/(1-\beta)}\,\mathbb{V}_{T}^{2/\alpha}\,T\ \lesssim\ J^{\,1+\frac{2}{\alpha q}}.$
Rearranging yields (211). Finally, the quantity that drives the regret separation in the information-limited regime is $\mu a^{2}T$. Using (209), plugging $J\asymp J_{\star}$ from (211), and simplifying exponents yields we get
$\inf_{\pi\in\Pi_{\beta}}\sup_{G:\,\mathrm{GVar}_{p,q}(G)\leq\mathbb{V}_{T}}\mathcal{R}_{T}^{\pi}(G)\;\gtrsim\;(1-\beta)^{-2/(\alpha q+2)}\sigma^{4/(\alpha q+2)}\mu^{(\alpha q-2q-2)/(\alpha q+2)}\mathbb{V}_{T}^{2q/(\alpha q+2)}T^{\alpha q/(\alpha q+2)}.$
E.6 Obtaining a minimax lower bound in the inertia-limited regime
Our information-theoretic construction (packing + Fano/KL) yields a lower bound that is driven by statistical indistinguishability of environments and therefore captures the dependence on the noise level $\sigma$ and the variation budget $\mathbb{V}_{T}$. However, it does not by itself explain the empirically dominant failure mode of momentum under drift: inertia. To isolate this mechanism, we analyze (SGDM) (Heavy-Ball) on the simplest strongly convex and smooth objective, $\phi_{u}(z)=\frac{\mu}{2}(z-ua)^{2}$, where the minimizer jumps between $\pm a$. We prove the momentum-specific lower bound by explicitly analyzing (SGDM) on a single block switch and showing it takes $\tau_{\beta}:=\Omega(\kappa/(1-\beta))$ steps to reduce the tracking error by a constant factor under a step size restriction. Crucially, this part does not rely on hiding information; the learner may instantly know the new function. The lower bound comes from the algorithmic constraint of (SGDM) under stability. We will focus on the Polyak Heavy-Ball method of momentum ($\beta_{1}=0,\beta_{2}=\beta$). However a similar analysis can be carried out for Nesterov and yield a analogous result.
Proposition E.1 ((SGDM) on a 1D quadratic).
Fix $u\in\{\pm 1\}$ and consider the one-dimensional $\mu$–strongly $\mu$-smooth convex quadratic $\phi_{u}(z)\;:=\;\frac{\mu}{2}(z-ua)^{2},\;\;z\in\mathbb{R}$. Let Heavy-Ball (SGDM) (Polyak momentum) with step size $\gamma>0$ and momentum $\beta\in[0,1)$ evolve as
$z_{t+1}=z_{t}+\beta(z_{t}-z_{t-1})-\gamma\big(\mu(z_{t}-ua)+\eta_{t}\big),\qquad\eta_{t}\sim\mathcal{N}(0,\sigma^{2})\ \text{i.i.d.}$ (213)
$\binom{\mathbb{E}[e_{t+1}]}{\mathbb{E}[e_{t}]}\;=\;\underbrace{\begin{pmatrix}1+\beta-\gamma\mu&-\beta\\ 1&0\end{pmatrix}}_{\stackrel{{\scriptstyle\Delta}}{{=}}\;A}\binom{\mathbb{E}[e_{t}]}{\mathbb{E}[e_{t-1}]}$, (214)
$0<\gamma\mu\leq\min\Big\{(1-\sqrt{\beta})^{2},\ \frac{1-\beta}{4}\Big\},$ (215)
$|\lambda_{\max}|\;\geq\;1-c\,\frac{\gamma\mu}{1-\beta}.$ (216)
$|\mathbb{E}[e_{t}]|\;\geq\;C^{-1}|\lambda_{\max}|^{\,t}\,a\;\geq\;C^{-1}\Big(1-c\,\frac{\gamma\mu}{1-\beta}\Big)^{t}a.$ (217)
In particular, defining the response time
$\tau_{\beta}\;:=\;\min\Big\{t\geq 0:\ \Big(1-c\,\frac{\gamma\mu}{1-\beta}\Big)^{t}\leq\frac{1}{2}\Big\},$ (218)
we have the scaling
$\tau_{\beta}\ \asymp\ \frac{1-\beta}{\gamma\mu}.$ (219)
$\mathbb{E}\big[\phi_{u}(z_{t})-\phi_{u}(ua)\big]=\frac{\mu}{2}\mathbb{E}[e_{t}^{2}]\;\geq\;\frac{\mu}{2}\big(\mathbb{E}[e_{t}]\big)^{2}\;\gtrsim\;\mu a^{2}\Big(1-c\,\frac{\gamma\mu}{1-\beta}\Big)^{2t}.$ (220)
Proof of Proposition˜E.1.
Subtracting $ua$ from (213) gives
$e_{t+1}=(1+\beta-\gamma\mu)e_{t}-\beta e_{t-1}-\gamma\eta_{t}.$ (221)
which is equivalently the two-dimensional linear system (214) with matrix $A$ as stated. The eigenvalues of $A$ are the roots of
$\lambda^{2}-(1+\beta-\gamma\mu)\lambda+\beta=0.$ (222)
Under (215), the discriminant is nonnegative so $\lambda_{\pm}\in\mathbb{R}$. Moreover, $\lambda_{\pm}\in(0,1)$ in this regime. Let $\lambda_{\max}:=\max\{\lambda_{+},\lambda_{-}\}$. To lower bound $\lambda_{\max}$, write $\delta:=1-\beta\in(0,1]$ and $\varepsilon:=\gamma\mu$. Then
$\lambda_{\max}=1-\frac{\delta+\varepsilon}{2}+\frac{1}{2}\sqrt{(\delta+\varepsilon)^{2}-4\varepsilon}.$
Let $a:=\delta+\varepsilon$ and $b:=4\varepsilon$. Under (215) the discriminant is nonnegative, i.e. $0\leq b\leq a^{2}$, so we may use $\sqrt{a^{2}-b}\geq a-\frac{b}{a}$ to obtain
$\lambda_{\max}\;\geq\;1-\frac{a}{2}+\frac{1}{2}\Big(a-\frac{b}{a}\Big)=1-\frac{b}{2a}=1-\frac{2\varepsilon}{\delta+\varepsilon}\;\geq\;1-\frac{2\varepsilon}{\delta}=1-2\,\frac{\gamma\mu}{1-\beta}$
which is (216) with a universal constant $c=2$. Since $A$ is diagonalizable with eigenvalues in $(0,1)$, the solution of (214) can be written as a linear combination of $\lambda_{+}^{t}$ and $\lambda_{-}^{t}$. For the symmetric initialization $\mathbb{E}[e_{0}]=\mathbb{E}[e_{-1}]=a$, the dominant mode is aligned with $\lambda_{\max}$, hence there exists a constant $C\geq 1$ (depending only on the conditioning of the eigenbasis, and thus universal in the regime (215)) such that (217) holds. The response time bound (219) follows immediately from the definition (218) and the standard approximation $\log(1-x)\asymp-x$ for $x\in(0,1/2)$, applied with $x=c\,\gamma\mu/(1-\beta)$. Finally, since $\phi_{u}(z_{t})-\phi_{u}(ua)=\frac{\mu}{2}e_{t}^{2}$ and $\mathbb{E}[e_{t}^{2}]\geq(\mathbb{E}[e_{t}])^{2}$, plugging (217) into this inequality yields (220). ∎
We now use this as the block length $\Delta_{T}\asymp\tau_{\beta}$. Since in our $J$-block construction, the minimizer is constant within a block and flips between blocks. If the blocks are much longer than $\tau_{\beta}$, (SGDM) has time to settle near the new minimizer after each switch. Then the regret per block is mostly transient and doesn’t accumulate strongly. On the other hand, if the block size is much shorter than $\tau_{\beta}$, (SGDM) will never be able to catch up. Taking $\Delta_{T}\asymp\tau_{\beta}$ will give us a constant per-round suboptimality throughout the block (up to constants), hence resulting in $\Omega(\mu a^{2}\Delta_{T})$ regret contribution per block. We formalize this in the following result:
Theorem E.1 (From response time to regret under block switching).
Consider the $J$-block construction of nonstationary losses $G^{u}_{1:T}$ over $[T]$ with blocks $B_{1},\dots,B_{J}$ of lengths $|B_{j}|\in\{\Delta_{T},\Delta_{T}+1\}$, $\Delta_{T}=\lfloor T/J\rfloor$, and $G_{t}^{u}\equiv g_{u_{j}}$ for $t\in B_{j}$, where $u\in\{\pm 1\}^{J}$. Assume the base losses are the translated quadratics $\phi_{\pm}$ above (embedded along $e_{1}$ if $d>1$), so that within each block the unique minimizer is $x^{\star}_{u_{j}}=u_{j}a$. Run Heavy-Ball (SGDM) with parameters $(\gamma,\beta)$ satisfying the stability cap
$\gamma\ \leq\ c_{0}\,\frac{(1-\beta)^{2}}{L},$ (223)
for a universal constant $c_{0}>0$. Let $\tau_{\beta}$ be the response time from Proposition˜E.1, so in particular
$\tau_{\beta}\ \gtrsim\ \frac{L}{\mu(1-\beta)}.$ (224)
Choose the block length on the order of the response time,
$\Delta_{T}\ \asymp\ \tau_{\beta},\qquad\text{equivalently}\qquad J\ \asymp\ \frac{T}{\tau_{\beta}}.$ (225)
Then there exists a universal constant $c>0$ such that for any $u\in\{\pm 1\}^{J}$ with a sign flip at each block boundary (e.g., $u_{j+1}=-u_{j}$), the expected dynamic regret of Heavy-Ball (SGDM) satisfies
$\mathcal{R}_{T}^{\pi_{\mathrm{HB}}}(G^{u})\;\geq\;c\,\mu a^{2}\,T.$ (226)
Moreover, if the block family is tuned to satisfy the gradient-variation budget $\mathrm{GVar}_{p,q}(G^{u}_{1:T})\leq\mathbb{V}_{T}$ via Lemma˜E.10, i.e.
[…]
then substituting $J\asymp T/\tau_{\beta}$ into (226) yields
$\mathcal{R}_{T}^{\pi_{\mathrm{HB}}}(G^{u})\;\gtrsim\;\mu^{1-2/\alpha}\,\mathbb{V}_{T}^{2/\alpha}\,\tau_{\beta}^{2/(\alpha q)}\,T^{1-(2/\alpha q)}.$ (228)
Finally, using (224) gives the explicit inertia-dependent lower bound
$\mathcal{R}_{T}^{\pi_{\mathrm{HB}}}(G^{u})\;\gtrsim\;\mu^{1-2/\alpha}\,\mathbb{V}_{T}^{2/\alpha}\,\Big(\frac{L}{\mu(1-\beta)}\Big)^{2/(\alpha q)}\,T^{1-(2/\alpha q)}.$ (229)
Proof of Theorem˜E.1.
Step 1: A flip induces an $\Omega(a)$ error at the start of the next block.
$e_{t}:=x_{t}-x^{\star}_{j+1},\qquad t\in B_{j+1},$
$|e_{t_{0}-1}|=\big|x_{t_{0}-1}-x^{\star}_{j+1}\big|\geq|x^{\star}_{j}-x^{\star}_{j+1}|-\big|x_{t_{0}-1}-x^{\star}_{j}\big|=2a-\big|x_{t_{0}-1}-x^{\star}_{j}\big|.$ (230)
Thus, whenever the iterate has achieved even a moderate accuracy on the previous block, say $|x_{t_{0}-1}-x^{\star}_{j}|\leq a$, we obtain $|e_{t_{0}-1}|\geq a$. In particular, for the alternating-sign choice in the theorem (a flip at every boundary), either the algorithm fails to track the previous minimizer by time $t_{0}-1$—which already incurs regret on block $B_{j}$—or else it necessarily starts block $B_{j+1}$ with initial error at least $a$ relative to the new minimizer. This is what allows us to lower bound regret block-by-block. To make this quantitative in a way compatible with the Heavy-Ball state, define the extended state
$s_{t}:=\begin{bmatrix}e_{t}\\ e_{t-1}\end{bmatrix}$.
$\|s_{t_{0}-1}\|_{2}\;\geq\;c_{\mathrm{init}}\,a.$ (231)
We will use (231) as the initial condition for the lag argument in Step 2.
Step 2: Over $\Theta(\tau_{\beta})$ steps the mean error cannot shrink by more than a constant factor.
On block $B_{j+1}$ the loss is the fixed quadratic $g_{u_{j+1}}(x)=\frac{\mu}{2}(x-x^{\star}_{j+1})^{2}$, so with $e_{t}:=x_{t}-x^{\star}_{j+1}$ the Heavy-Ball recursion gives
$e_{t+1}=(1+\beta-\gamma\mu)e_{t}-\beta e_{t-1}-\gamma\eta_{t},\qquad t\in B_{j+1}.$
Let $\bar{e}_{t}:=\mathbb{E}[e_{t}]$ and $\bar{s}_{t}:=\begin{bmatrix}\bar{e}_{t}\\ \bar{e}_{t-1}\end{bmatrix}$. Taking expectations yields the homogeneous linear system
$\bar{s}_{t+1}=A\,\bar{s}_{t},\qquad A=\begin{bmatrix}1+\beta-\gamma\mu&-\beta\\ 1&0\end{bmatrix}$. (232)
Fix a flipped boundary between $B_{j}$ and $B_{j+1}$, and let $t_{0}$ be the first index of $B_{j+1}$. By Step 1, either the previous block already contributes $\gtrsim\mu a^{2}\Delta_{T}$ regret (and we are done for that block), or else the algorithm is at least moderately aligned with the previous minimizer in the sense that $|\mathbb{E}[x_{t_{0}-1}]-x^{\star}_{j}|\leq a$ and $|\mathbb{E}[x_{t_{0}-2}]-x^{\star}_{j}|\leq a$. In this latter case, using $x^{\star}_{j+1}=-x^{\star}_{j}$ and the reverse triangle inequality for scalars gives
$|\bar{e}_{t_{0}-1}|=|\mathbb{E}[x_{t_{0}-1}]-x^{\star}_{j+1}|\geq|x^{\star}_{j}-x^{\star}_{j+1}|-|\mathbb{E}[x_{t_{0}-1}]-x^{\star}_{j}|\geq 2a-a=a,$
and similarly $|\bar{e}_{t_{0}-2}|\geq a$. Thus $\bar{s}_{t_{0}-1}$ satisfies $|\bar{e}_{t_{0}-1}|,|\bar{e}_{t_{0}-2}|\geq a$.
Now apply Proposition˜E.1 to the shifted mean trajectory on this block (i.e., re-index time so that $t_{0}-1$ plays the role of 0). Using (217) and the definition (218), we obtain that there exists a universal constant $c_{\mathrm{lag}}\in(0,1)$ such that for all $k=0,1,\dots,\lfloor\tau_{\beta}\rfloor$,
$|\bar{e}_{t_{0}+k}|\;\geq\;c_{\mathrm{lag}}\,a.$ (233)
Consequently, Jensen implies that on the same window
$\mathbb{E}[e_{t_{0}+k}^{2}]\ \geq\ (\mathbb{E}[e_{t_{0}+k}])^{2}\ \geq\ c_{\mathrm{lag}}^{\,2}\,a^{2}.$ (234)
For the quadratic loss on block $B_{j+1}$,
$g_{u_{j+1}}(x_{t})-g_{u_{j+1}}(x^{\star}_{j+1})=\frac{\mu}{2}\,e_{t}^{2}\;\geq\;\frac{\mu}{2}\,\Big(\,|e_{t}|\,\Big)^{2}.$ (235)
By (234), for each $k\leq\lfloor c_{\mathrm{lag}}\tau_{\beta}\rfloor$ we have $|e_{t_{0}+k}|\gtrsim a$, hence the per-round regret on that window is
$g_{t_{0}+k}(x_{t_{0}+k})-g_{t_{0}+k}(x^{\star}_{t_{0}+k})\;=\;g_{u_{j+1}}(x_{t_{0}+k})-g_{u_{j+1}}(x^{\star}_{j+1})\;\gtrsim\;\mu a^{2}.$ (236)
Summing (236) over $k=0,\dots,\lfloor c_{\mathrm{lag}}\tau_{\beta}\rfloor$ gives an expected regret contribution of order $\mu a^{2}\tau_{\beta}$ per flipped block:
$\sum_{t\in B_{j+1}}\big(g_{t}(x_{t})-g_{t}(x_{t}^{\star})\big)\;\geq\;\sum_{k=0}^{\lfloor c_{\mathrm{lag}}\tau_{\beta}\rfloor}\big(g_{t_{0}+k}(x_{t_{0}+k})-g_{t_{0}+k}(x^{\star}_{t_{0}+k})\big)\;\gtrsim\;\mu a^{2}\,\tau_{\beta}.$ (237)
With the block-length choice $\Delta_{T}\asymp\tau_{\beta}$ in (225), this becomes $\gtrsim\mu a^{2}\,\Delta_{T}$ per block. Since the construction flips at every boundary, a constant fraction of the $J$ blocks contribute this amount, and therefore
$\mathcal{R}_{T}^{\pi_{\mathrm{HB}}}(F^{u})=\sum_{t=1}^{T}\big(g_{t}(x_{t})-g_{t}(x_{t}^{\star})\big)\;\gtrsim\;J\cdot\mu a^{2}\,\Delta_{T}\;\asymp\;\mu a^{2}\,T,$
which proves (226).
Under the $J$-block construction, Lemma˜E.10 gives $\mathrm{GVar}_{p,q}(G^{u}_{1:T})\lesssim\mu a^{\alpha}J^{1/q}$. Enforcing $\mathrm{GVar}_{p,q}(G^{u}_{1:T})\leq\mathbb{V}_{T}$ yields (227) (absorbing constants). Plugging (227) into (226) gives
$\mathcal{R}_{T}^{\pi_{\mathrm{HB}}}(G^{u})\ \gtrsim\ \mu T\Big(\frac{\mathbb{V}_{T}}{\mu J^{1/q}}\Big)^{\!2/\alpha}=\mu^{1-2/\alpha}\,\mathbb{V}_{T}^{2/\alpha}\,T\,J^{-2/(\alpha q)}.$
Using $J\asymp T/\tau_{\beta}$ from (225) yields (228). Finally, the stability cap (223) implies $\gamma\lesssim(1-\beta)^{2}/L$, and combining this with $\tau_{\beta}\asymp(1-\beta)/(\gamma\mu)$ from Proposition˜E.1 yields $\tau_{\beta}\gtrsim L/(\mu(1-\beta))$, i.e. (224). Substituting this lower bound into (228) gives (229). ∎
Since both the statistical and inertia constructions produce $\mu$-strongly convex, $\mu$-smooth functions satisfying $\mathrm{GVar}_{p,q}\leq\mathbb{V}_{T}$, the restricted minimax regret over $\Pi_{\beta}$ is at least the maximum of the two lower bounds. This completes the proof of the result.
Appendix F Additional experimental details and results
We provide complete details for reproducibility. All experiments compare SGD, Heavy-Ball (HB), and Nesterov (NAG) under identical drift conditions, with step sizes selected by a fixed tuning protocol described below.
F.1 General setup
Step size constraints and drift model.
Our results (Theorem˜3.3 and Theorem˜3.6) assumes the sufficient stability condition $\gamma\leq\mu(1-\beta)^{2}/(4L^{2})$ for HB/NAG to obtain clean finite-time bounds. In the main experimental setup, we do not enforce this conservative cap: we evaluate all methods at a standard set of base step sizes $\gamma$ reported in the tables (and used uniformly across methods within each task/regime) to test whether the drift-induced inertia phenomenon persists beyond the theoretically analyzed regime. For reference, when we choose step sizes within the stability region, we observe the same qualitative behavior as (Yuan et al., 2016). Across all tasks, we evolve the population minimizer $\bm{\theta}_{t}^{\star}\in\mathbb{R}^{d}$ via a normalized random walk